RHIC pp Set-Up: LINAC Failure lost 2 hours.
Blue stochastic cooling test, moved the storage and Landau RF phases at store for Blue.
Yellow RF had wrong gamma_t value in DSP code. Necessitated half bucket timing shift at store (17ns) and wrong chromaticity. DX BPMs at store still need to be adjusted.
Non-linear IR correction needs work.
Overnight collisions with 36x36 1e11. PHENIX ZDC of 2kHz, 50kHz in BBC.
Today, continue Nonlinear IR correction in Blue (Fulvia), Nonlinear chromaticity work (Yun), Feedback ramps and decoupling (Yun), Ramp tuning for polarization, keep tune swing. Overnight collisions at 1.1e11 56x56. With pp83 ramp, not pp83ro. I.e. continue with vertical polarization until PHENIX gets a good run at 40% polarization.
CNI measurements every hour. Get CNI into CDEV.
Continue Machine Development over the weekend.
STAR needs a half hour access to fix a trigger. PHENIX will take advantage of the access to replace a LV PS. PP2PP will also get work done inside. This access will take place as soon as the morning meeting is over.
STAR reports a twenty second ripple in the background from Yellow.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |